A highway safety study found that in 84% of all accidents, the driver was wearing a seatbelt. Accident reports indicate that 90% of those drivers escaped from serious injury (hospitalized or death) but only 60% of the non-belted drivers were so lucky.  

What is the probability that a seriously injured driver was not wearing a seatbelt.
Answer #1
P(seriosly injured)=P(not wearing)*P(seriosly injured|not wearing)+P(wearing)*P(seriosly injured|wearing)
=(1-0.84)*(1-0.4)+0.84*(1-0.1)=0.148
therefore P(not wearing|seriosly injured)=P(not wearing)*P(seriosly injured|not wearing)/P(seriosly injured)=(1-0.84)*(1-0.6)/0.148 =0.4324